# CAIMAN: Causal Action Influence Detection for Sample-efficient Loco-manipulation

### v2

@@ Abstract
Enabling legged robots to perform non-prehensile loco-manipulation is crucial for enhancing their versatility. Learning behaviors such as whole-body object pushing often requires sophisticated planning strategies or extensive task-specific reward shaping, especially in unstructured environments. In this work, we present CAIMAN, a practical reinforcement learning framework that encourages the agent to gain control over other entities in the environment. CAIMAN leverages causal action influence as an intrinsic motivation objective, allowing legged robots to efficiently acquire object pushing skills even under sparse task rewards. We employ a hierarchical control strategy, combining a low-level locomotion module with a high-level policy that generates task-relevant velocity commands and is trained to maximize the intrinsic reward. To estimate causal action influence, we learn the dynamics of the environment by integrating a kinematic prior with data collected during training. We empirically demonstrate CAIMAN’s superior sample efficiency and adaptability to diverse scenarios in simulation, as well as its successful transfer to real-world systems without further fine-tuning.
Keywords: Reinforcement Learning, Loco-manipulation, Intrinsic Motivation
1 Introduction
@@
Modern legged robots continue to impress with their diverse capabilities, from traversing challenging terrains [1, 2, 3] to executing agile maneuvers such as backflips and parkour [4, 5, 6]. However, enabling these systems to interact with their environments remains an open research challenge [7, 8, 9]. A common approach is to equip legged robots with external manipulators for prehensile manipulation [10, 11, 12, 13]. Although effective, these methods are often constrained by the size and weight of objects they can successfully manipulate. In contrast, non-prehensile manipulation demands whole-body coordination and remains a non-trivial task.
@@
Traditional approaches for whole-body manipulation explicitly model robot-object interactions, often relying on complex planning and optimization for coordinating motion [14]. These methods require accurate models of both the robot and the environment, making them less scalable to high-dimensional systems with complex, stochastic dynamics, and often limiting contact to predefined regions [15, 16]. Learning-based methods offer a more scalable alternative, improving computational efficiency and reducing dependence on precise object estimation [17, 18, 19]. However, these methods frequently rely on tedious reward shaping or learning curriculums to guide exploration and foster meaningful behavior. Furthermore, the large exploration space of loco-manipulation tasks often necessitates special treatment, such as learning from behavioral priors  [20, 21, 22] or task-agnostic explorative rewards [23, 24], to encourage meaningful engagement with the environment. Prior works have also leveraged a hierarchical structure to decouple locomotion and high-level planning [25]. Training the high-level policy, however, still requires elaborate reward design and shaping.
@@
To alleviate these challenges, we introduce CAIMAN, a framework for training non-prehensile object pushing skills in legged robots. CAIMAN employs a hierarchical control structure that decouples high-level planning from low-level locomotion. We train robust locomotion using an existing pipeline, and we learn the high-level policy with a simple yet effective reward structure consisting of only three terms: a sparse task reward, an action regularizer, and an intrinsically motivated explorative reward. The sparse task reward provides a general signal for various pushing tasks, including ones requiring navigation through obstacles. The high-level policy is incentivized to explore and gain control over the environment via an exploratory reward based on Causal Action Influence (CAI) [26], which quantifies how much influence the agent has on the states of other entities and is computed based on environment dynamics. Instead of learning the dynamics from scratch, CAIMAN utilizes a simple pre-defined kinematic prior, and combines it with learned residual dynamics that compensate for physical interactions beyond the prior model. Using only the sparse task reward, we show that CAIMAN achieves strong learning performance and superior sample efficiency against other baselines, especially in complex scenarios with obstacles. We also demonstrate that the learned residual accurately captures complex robot-object interactions that are not modeled by the kinematic prior. Finally, we successfully transfer the learned policy to a real quadruped robot, enabling it to perform whole-body object pushing in obstacle environments.
@@
To summarize, our contribution is three-fold: 1) a general hierarchical framework for learning whole-body object pushing with legged robots in various scenarios, including navigating through obstacles; 2) an intrinsically motivated reward based on causal influence calculated from a combination of a kinematics prior and learned residual dynamics; 3) Successful hardware validation on a real-world quadruped robot.
2 Related Work
2.1 Loco-manipulation on Legged Systems
@@
The integration of locomotion and manipulation has gained significant attention as a promising and application-oriented research field for legged robots. Traditional model-based methods that rely on accurate representations of both robot and object to optimize trajectories [11, 14, 16, 27] have shown success in tasks such as box carrying [10, 12], but often require accurate state estimation [28, 29, 30] and struggle with scalability due to the complexity of contact modes [31, 32]. Reinforcement learning (RL) is a promising alternative that avoids explicit modeling and has been successfully applied to tasks such as soccer dribbling [33, 34], button pushing [18, 35], and door opening [19, 23, 36]. However, effective exploration may be difficult to achieve due to the large search space of robot-object interactions [17, 37]. To address exploration challenges, researchers have proposed using behavior prior [38, 39] or task-agnostic explorative rewards [23, 24] as mechanisms to guide learning. Furthermore, hierarchical frameworks [40, 41] are utilized to decompose tasks into high-level planning and low-level control, enabling effective behavior across various scenarios [42, 43].
@@
In accordance with previous works, we also implement a hierarchical control framework for whole-body object pushing, decoupling high-level velocity planning from low-level locomotion, similarly to Jeon et al. [25]. We aim to utilize a simple reward structure, as opposed to one that requires sophisticated design effort as seen in [25], to achieve sample-efficient high-level policy training, even under complex scenarios such as object manipulation through obstacles.
2.2 Intrinsically Motivated Reinforcement Learning
@@
Intrinsic motivation (IM) [44] plays a vital role in reinforcement learning, especially when extrinsic rewards are sparse or difficult to design. Core IM mechanisms include curiosity [45], learning progress [46], and empowerment [47], each promoting exploration and skill acquisition. Curiosity, often measured via prediction errors in learned world models [48, 49, 50], rewards agents for encountering novel states, and has been previously applied for learning loco-manipulation skills [23, 24]. Learning progress encourages agents to focus on regions in the state space with rapid improvement, supporting curriculum learning and adaptive exploration [51, 52].
@@
Our work builds on empowerment, an information-theoretic quantity defined as the channel capacity between the agent’s actions and its sensory observations [47, 53, 54]. Recent work has connected IM to causal reasoning, aiming to improve sample efficiency and interpretability [55, 56, 57, 58, 59]. To encourage effective exploration, we employ causal action influence (CAI) [26], a measure of an agent’s ability to influence its environment and a conceptual lower bound on empowerment.
3 Preliminaries
(a) No influence of $A$ on $S^{\prime}_{2}$
(b) Influence of $A$ on $S^{\prime}_{1}$ and $S^{\prime}_{2}$
@@ 3 Preliminaries
Figure 1: Illustration of the LCM (left) for two different environment situations $S=s$ (right) in the loco-manipulation task. The LCM captures the transition from $S,A$ to $S^{\prime}$, factorized into state components. While the global SCM is fully connected (dashed and continuous lines), the LCM $\mathcal{G}_{S=s}$ (continuous lines) is causally minimal. We are interested in detecting the presence of continuous orange arrows in the LCM, i.e. the influence of the action $A$ on next states $S^{\prime}$.
@@
We model decision-making in a dynamic environment as a Markov Decision Process (MDP) [60], defined by the tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P$ the transition kernel, $R$ the reward function, and $\gamma$ the discount factor. Following the principle of independent causal mechanisms [61], we assume that the world consists of interacting but independent entities. This induces a state space factorization $\mathcal{S}=\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{N}$ for $N$ entities, where each factor $\mathcal{S}_{i}$ represents the state of entity $i$. An MDP coupled with a policy $\pi:\mathcal{S}\mapsto\mathcal{A}$ induces a Structural Causal Model (SCM) [62] describing the resulting trajectory distribution.
Definition 1 (Structural Causal Model [62]).
@@
An SCM is a tuple $(\mathcal{U},\mathcal{V},F,P^{u})$, where $\mathcal{U}$ are set of exogenous variables (e.g., latent randomness) sampled from $P^{u}$, $\mathcal{V}$ are set of observed variables (e.g., states, actions, rewards), $F$ is the set of structural functions capturing the causal relations, such that functions $f_{V}:\mathrm{Pa}(V)\times\mathcal{U}\rightarrow V$, with $\mathrm{Pa}(V)\subset\mathcal{V}$ denoting the set of parents of $V$, determine the value of endogenous variables $V$ for each $V\in\mathcal{V}$.
@@
SCMs are typically visualized as directed acyclic graphs, where nodes represent variables and edges indicate causal relations, as shown in Figure 1. In our case, the SCM captures one-step transitions with variables $\mathcal{V}={S_{1},\ldots,S_{N},A,S_{1}^{\prime},\ldots,S_{N}^{\prime}}$. Due to the Markov property and flow of time, causal dependencies exist only from $(S,A)$ to $S^{\prime}$. While most entity pairs may theoretically interact (i.e., $S_{i}/A\rightarrow S^{\prime}_{j}$ for most $i,j$), interactions often become sparse when we observe a specific state configuration. We use the Local Causal Model (LCM) to capture such context-specific structure [63].
Definition 2 (Local Causal Model [63]).
@@
In this work, we build on the insight that encouraging agents to gain control over their environment facilitates learning in loco-manipulation tasks. We do so by driving the agent toward states where it can influence other entities. We make use of a principled explicit measure of local influence, the Causal Action Influence (CAI) [26, 64] measure. CAI is a state-dependent measure of control that assesses whether an agent, through its actions, can affect other entities. Graphically, it corresponds to predicting the existence of an edge $A\rightarrow S^{\prime}_{j}$ in the LCM $\mathcal{G}_{do(S=s)}$. This influence is measured with point-wise conditional mutual information $I(S_{j}^{\prime};A\mid S=s)$.
4 Method
Figure 2: CAIMAN framework: The high-level (HL) policy generates desired base velocity commands, which are translated into joint commands by a low-level (LL) policy. We utilize a simple kinematic prior and learned residual dynamics to model the robot-object interaction in the environment while providing a CAI-based explorative bonus along with the sparse task reward.
@@ 4 Method
The hierarchical framework adopted by CAIMAN is presented in Figure 2. It consists of a high-level (HL) policy that generates desired base velocity commands and a low-level policy that translates velocity commands into joint-level actions. In this work, we introduce a novel, sample-efficient approach for learning the high-level policy, while building upon an existing pipeline [65] to train robust low-level policies.
4.1 Low-level Locomotion Policy

### v3

@@ Abstract
Enabling legged robots to perform non-prehensile loco-manipulation is crucial for enhancing their versatility. However, learning behaviors such as whole-body object pushing often necessitates sophisticated planning strategies or extensive task-specific reward shaping. In this work, we present CAIMAN, a practical reinforcement learning framework that encourages the agent to gain control over other entities in the environment. CAIMAN leverages causal action influence as an intrinsic motivation objective, allowing legged robots to efficiently acquire object pushing skills even under sparse task rewards. We employ a hierarchical control strategy, combining a low-level locomotion module with a high-level policy that generates task-relevant velocity commands and is trained to maximize the intrinsic reward. To estimate causal action influence, we learn the dynamics of the environment by integrating a kinematic prior with data collected during training. We empirically demonstrate CAIMAN’s superior sample efficiency and adaptability to diverse scenarios in simulation, as well as its successful transfer to real-world systems without further fine-tuning. A video demo is available at https://www.youtube.com/watch?v=dNyvT04Cqaw.
I INTRODUCTION
@@
Modern day legged robots showcase impressive versatility, from traversing challenging terrains [miki2022learning, choi2023learning] to executing agile maneuvers such as backflips and parkour [katz2019mini, cheng2024extreme]. Yet as expectations for autonomy grow, enabling these systems to physically interact with their environments remains a central research problem [ha2024learning, tang2024deep, gu2025humanoid]. A common strategy to enhance manipulation capabilities is to equip legged robots with external manipulators for prehensile tasks [bellicoso2019alma, zimmermann2021go, sleiman2021unified], but such methods are inherently constrained by object size and payload. Harnessing whole-body motion for non-prehensile manipulation offers a promising alternative, but remains a non-trivial challenge.
@@
Traditional approaches for whole-body manipulation explicitly model robot-object interactions, often relying on complex planning and optimization for coordinating motion [murooka2015whole]. These methods require accurate models of both the robot and the environment, which restricts their scalability for high-dimensional systems with complex, stochastic dynamics and limits contact points to predefined regions [farnioli2016toward, rigo2023contact]. Learning-based methods offer a more scalable alternative, improving computational efficiency and reducing dependence on precise object estimation [shi2021circus, cheng2023legs, arm2024pedipulate]. However, these methods frequently rely on tedious reward shaping or learning curriculums to guide exploration and foster meaningful behavior. Furthermore, the large exploration space of loco-manipulation tasks often necessitates special treatment, such as learning from behavioral priors  [ha2024umi, urpi2023efficient] or task-agnostic explorative rewards [schwarke2023curiosity, zhang2024wococo], to encourage meaningful engagement with the environment.
Figure 1: Robot loco-manipulation in the real world enabled by CAIMAN. The quadruped maneuvers a box around an obstacle to reach a target position.
@@
To alleviate these challenges, we introduce CAIMAN, a framework for training non-prehensile object pushing skills in legged robots. CAIMAN employs a hierarchical control structure that decouples high-level planning from low-level locomotion. We train robust locomotion using an existing pipeline, and we learn the high-level policy with a simple yet effective reward structure consisting of only three terms: a sparse task reward, an action regularizer, and an intrinsically motivated explorative reward. The sparse task reward provides a general signal for various pushing tasks, where the robot is only rewarded for completing the task successfully. The exploratory reward incentivizes the robot to explore and gain control over the environment via Causal Action Influence (CAI) [seitzer2021causal], which is a measure quantifying how much influence an agent has on the states of other entities and is computed based on environment dynamics. Instead of learning the dynamics from scratch, CAIMAN combines a simple predefined kinematic prior with learned residual dynamics that compensate for physical interactions beyond the prior, allowing accurate dynamics to be learned efficiently. We show that in a sparse task reward setting, CAIMAN achieves strong learning performance and superior sample efficiency compared to other baselines, including within complex, obstacle-laden scenarios. We also demonstrate that the learned residual accurately captures complex robot-object interactions that are not modeled by the kinematic prior. Finally, we successfully transfer the learned policy to a real quadruped robot, enabling it to perform whole-body object pushing in real-world settings.
@@
To summarize, our contribution is three-fold: 1) a general hierarchical framework for learning whole-body object pushing with legged robots in various scenarios, including navigation through obstacles; 2) an intrinsically motivated reward based on causal influence calculated from a combination of a kinematics prior and learned residual dynamics; 3) successful hardware validation on a real-world quadruped robot.
II RELATED WORK
II-A Loco-manipulation on Legged Systems
@@
The integration of locomotion and manipulation has gained significant attention as a promising and application-oriented research field for legged robots. Model-based methods that rely on accurate representations of both robot and object to optimize trajectories [zimmermann2021go, murooka2015whole, rigo2023contact] have shown success in tasks such as box carrying [bellicoso2019alma, sleiman2021unified], but often require precise state estimation [lin2024locoman, schakkal2024dynamic] and struggle with scalability due to the complexity of contact modes [cheng2023enhancing]. Reinforcement learning (RL) is a viable alternative that avoids explicit modeling and has been successfully applied to diverse tasks including soccer dribbling [ji2023dribblebot, hu2024dexdribbler], button pushing [cheng2023legs, he2024learning], and door opening [arm2024pedipulate, schwarke2023curiosity]. However, achieving effective exploration remains difficult given the large search space of robot-object interactions [shi2021circus, fu2023deep]. To address exploration challenges, researchers have proposed using behavior prior [sleiman2024guided, liu2024opt2skill] or task-agnostic explorative rewards [schwarke2023curiosity, zhang2024wococo] as mechanisms to guide learning. In addition, hierarchical frameworks [rigo2024hierarchical, wang2024hypermotion] decompose tasks into high-level planning and low-level control, enabling effective behaviors across various scenarios [kumar2023cascaded].
@@
In accordance with previous works, we also implement a hierarchical control framework for whole-body object pushing, decoupling high-level velocity planning from low-level locomotion, similarly to Jeon et al. [jeon2023learning]. We aim to utilize a simple reward structure, as opposed to one that requires sophisticated design effort as seen in [jeon2023learning], to achieve sample-efficient high-level policy training.
II-B Intrinsically Motivated Reinforcement Learning
@@
Intrinsic motivation (IM) [rm2000intrinsic] plays a vital role in reinforcement learning, especially when extrinsic rewards are sparse or difficult to design. Core IM mechanisms include curiosity [schmidhuber1991possibility], learning progress [schmidhuber2010formal], and empowerment [klyubin2005empowerment], each promoting exploration and skill acquisition. Curiosity, often measured via prediction errors in learned world models [pathak2017curiosity, burda2019exploration], rewards agents for encountering novel states, and has been previously applied for learning loco-manipulation skills [schwarke2023curiosity, zhang2024wococo]. Learning progress encourages agents to focus on regions in the state space with rapid improvement, supporting curriculum learning and adaptive exploration [blaes2019control, colas2019curious].
@@
Our work builds on empowerment, an information-theoretic quantity defined as the channel capacity between the agent’s actions and its sensory observations [klyubin2005empowerment, mohamed2015variational]. Recent work has connected IM to causal reasoning, aiming to improve sample efficiency and interpretability [buesing2018woulda, sontakke2021causal]. To encourage effective exploration, we employ causal action influence (CAI) [seitzer2021causal], a measure of an agent’s ability to influence its environment and a conceptual lower bound on empowerment.
III PRELIMINARIES
(a) No influence of AA on $S^{\prime}_{2}$
(b) Influence of AA on $S^{\prime}_{1}$ and $S^{\prime}_{2}$
@@ III PRELIMINARIES
Figure 2: Illustration of the LCM (left) for two different environment situations $S=s$ (right) in the loco-manipulation task. The LCM captures the transition from $S,A$ to $S^{\prime}$, factorized into state components. The global SCM is fully connected (dashed and continuous lines), while the LCM $\mathcal{G}_{S=s}$ (continuous lines) is sparser. We are interested in detecting the presence of continuous orange arrows in the LCM, i.e. the influence of the action AA on next states $S^{\prime}$.
@@
We model decision-making in a dynamic environment as a Markov Decision Process (MDP) [sutton2018reinforcement], defined by the tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P$ the transition kernel, $R$ the reward function, and $\gamma$ the discount factor. Following the principle of independent causal mechanisms [peters2017elements], we assume that the world consists of interacting but independent entities. This induces a state space factorization $\mathcal{S}=\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{N}$ for $N$ entities, where each factor $\mathcal{S}_{i}$ represents the state of entity $i$. An MDP coupled with a policy $\pi:\mathcal{S}\mapsto\mathcal{A}$ induces a Structural Causal Model (SCM) [pearl2009causality] describing the resulting trajectory distribution.
Figure 3: CAIMAN framework: The high-level (HL) policy generates desired base velocity commands, which are translated into joint commands by a low-level ($L$) policy. We utilize a simple kinematic prior and learned residual dynamics to model the robot-object interaction in the environment while providing a CAI-based explorative bonus along with the sparse task reward.
Definition 1 (Structural Causal Model [pearl2009causality])
@@
An SCM is a tuple $(\mathcal{U},\mathcal{V},F,P^{u})$, where $\mathcal{U}$ is a set of exogenous variables (e.g., latent randomness) sampled from $P^{u}$, $\mathcal{V}$ is a set of observed variables (e.g., states, actions, rewards), and $F$ is the set of structural functions capturing the causal relations, such that functions $f_{V}:\mathrm{Pa}(V)\times\mathcal{U}\rightarrow V$, with $\mathrm{Pa}(V)\subset\mathcal{V}$ denoting the set of parents of $V$, determine the value of endogenous variables $V$ for each $V\in\mathcal{V}$.
@@
SCMs are typically visualized as directed acyclic graphs, where nodes represent variables and edges indicate causal relations, as shown in Figure 2. The SCM we consider captures one-step transitions with variables $\mathcal{V}={S_{1},\ldots,S_{N},A,S_{1}^{\prime},\ldots,S_{N}^{\prime}}$. Due to the Markov property and flow of time, causal dependencies exist only from $(S,A)$ to $S^{\prime}$. The global SCM is generally fully connected, encoding all possible, however unlikely, interactions between entities (i.e., $S_{i}/A\rightarrow S^{\prime}_{j}$ for many $i,j$) (see  Figure 2). However, most of these dependencies are inactive given a specific state configuration. For example, the robot can only influence an object if it is within a certain proximity to it. To capture such context-specific structure, we use the Local Causal Model (LCM) [pitis2020counterfactual].
Definition 2 (Local Causal Model [pitis2020counterfactual])
@@
Intuitively, the LCM is sparser than the global SCM, as it retains only the edges corresponding to entity influences that are active in the current context. This ensures that the graph reflects the agent’s local influences, rather than all potential ones.
@@
In this work, we build on the insight that agents can effectively learn loco-manipulation tasks when encouraged to gain control over their environment. We do so by driving the agent toward states where it can influence other entities, achieved through a principled explicit measure of local influence, the Causal Action Influence (CAI) [seitzer2021causal, urpicausal] measure. CAI is a state-dependent measure of control that assesses whether an agent’s actions can affect the states of other entities, corresponding graphically to the presence of an edge $A\rightarrow S^{\prime}_{j}$ in the LCM $\mathcal{G}_{do(S=s)}$. Formally, CAI is defined via point-wise conditional mutual information $I(S_{j}^{\prime};A\mid S=s)$, which equals zero when $S_{j}^{\prime}$ is independent of $A$ given $S=s$, i.e. $S_{j}^{\prime}\perp\!\!\!\perp A\mid S=s$. At state $S=s$, CAI is given by
$\begin{split}C^{j}(s)&:=I(S^{\prime}_{j};A\mid S=s)\\ &=\mathbb{E}_{a\sim\pi}\big[D_{KL}\big(P_{S^{\prime}_{j}\mid S=s,A=a}\;\big|\big|\;P_{S^{\prime}_{j}\mid S=s}\big)\big].\end{split}$ (1)
IV METHOD
@@
The hierarchical control and training framework adopted by CAIMAN is presented in Figure 3. It consists of a high-level (HL) policy that generates desired base velocity commands and a low-level policy that translates velocity commands into joint-level actions. In this work, we introduce a novel, sample-efficient approach for learning the high-level policy, while building upon an existing pipeline [rudin2022learning] to train robust low-level policies.
IV-A Low-level Locomotion Policy